# Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v3

@@ Abstract.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers. We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency and introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings.
@@
journalyear: 2025copyright: ccconference: The 16th ACM International Conference on Future and Sustainable Energy Systems; June 17–20, 2025; Rotterdam, Netherlandsbooktitle: The 16th ACM International Conference on Future and Sustainable Energy Systems (E-ENERGY ’25), June 17–20, 2025, Rotterdam, Netherlandsdoi: 10.1145/3679240.3734614isbn: 979-8-4007-1125-1/2025/06ccs: Social and professional topics Sustainabilityccs: Networks Cloud computing
@@ 2. Carbon-Aware QoR Adaptation
[…]
@@ 3. Online Approach
Based on these findings, we propose a multi-horizon optimization which periodically updates forecasts and determines an optimal (or close-to-optimal) deployment and load balancing for every interval $i$, described in Algorithm 1. We perform the optimization in two distinct steps:
@@
Algorithm 1 Multi-Horizon Optimization
1:for $\alpha\leftarrow 0$ to $I-1$ do
2:    # Long-term optimization
3:   if $\alpha\equiv 0\pmod{\tau}$ then
4:      $(R^{i},C^{i})_{i=\alpha}^{I-1}\leftarrow$ update forecasts
6:    # Short-term optimization
7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}}\sum_{i=\alpha}^{I-1}E^{i}$
8:    # Progress in time
@@
10:   $R^{\alpha},C^{\alpha},D^{\alpha},A^{\alpha}\leftarrow$ update past with observed reality
@@ 4. Experiments
[…]
@@
Carbon intensity traces.  We evaluate on ten globally distributed regions: California, US (CISO), Texas, US (ERCOT), New York, US (NYISO), Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), Sweden (SE), and Queensland, Australia (AU-QLD). All carbon intensity data was provided by ElectricityMaps (Maps, 2024). Our forecast methodology is explained in Appendix E.
@@ 4.2. Upper-Bound Potential
[…]
@@ 4.3. Potential under Realistic Conditions
Figure 4 shows the gap to the upper-bound performance. Across all experiments, savings under realistic conditions reached $82\pm 6$ of the upper-bound potential. Notably, even in regions with highly unpredictable request patterns (Random and Cell B, D and F), the online optimization does not show a significant performance drop.

### v4

@@ Abstract.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers. We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency. Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget. Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
@@
journalyear: 2025copyright: noneconference: The 16th ACM International Conference on Future and Sustainable Energy Systems; June 17–20, 2025; Rotterdam, Netherlandsbooktitle: The 16th ACM International Conference on Future and Sustainable Energy Systems (E-ENERGY ’25), June 17–20, 2025, Rotterdam, Netherlandsdoi: 10.1145/3679240.3734614isbn: 979-8-4007-1125-1/2025/06ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingThis manuscript is an extended version of our paper published at ACM e-Energy’25 (DOI: 10.1145/3679240.3734614). Compared to the published version, we (1) add a time-based vs. utilization-based power attribution perspective together with a proof that both yield equivalent provisioning decisions under mild assumptions (Appendix H), (2) extend the online approach with an automatic adaptation of $\text{QoR}_{\text{target}}$ to meet a fixed annual carbon budget (Sections 3.1 and 4.4).
@@ 2. Carbon-Aware QoR Adaptation
Problem setting.  Given an interval duration $\Delta$, we divide our optimization time window $[0,T]$ into $I=T/\Delta$ time intervals, indexed by $i$. For the remainder of this paper, we consider $T=1~\text{year}$ and $\Delta=1$ hour. We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$, and a tuple of two service quality tiers $\mathcal{Q}=(\text{Tier~1},\text{Tier~2})$. Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances. For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group. Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$. The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$). The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval. Appendix A shows an overview of all input and decision variables.
@@
Time-based vs. utilization-based power attribution.  The operational emissions in (2) require a choice of how to attribute power draw to a service and its resources. Two common modeling choices are: (i) utilization-based power attribution, where per-machine power depends on resource utilization, and (ii) time-based power attribution, where every active machine is assigned a (typically constant) attributed power draw per unit time.
Utilization-based models are a natural fit when modeling operator-side electricity usage (often reported as Scope 2), where metering reflects the actual energy consumed by servers. Time-based attribution is common in tenant-side carbon accounting (often reported as Scope 3) when emissions are allocated proportional to instance-hours. However, Scope 3 methodologies can also be utilization- or energy-based, and Scope 2 reporting can rely on attributed values depending on available measurements.
For utilization-based attribution, machine power draw can be modeled as utilization-dependent, e.g., with a concave power law,
(3) $p^{i}_{q}=p^{\text{idle}}+\bigl(p_{q}^{\text{max}}-p^{\text{idle}}\bigr)\cdot\bigl(\text{util}^{i}_{q}\bigr)^{n},\quad n\in(0,1],$
where, in our setting with one user group and one machine type, $\text{util}^{i}_{q}=a^{i}_{q}/(d^{i}_{q}\cdot k_{q})$. Under time-based attribution, $p^{i}_{q}$ is treated as utilization-independent (constant per active instance). We show that under mild concavity assumptions, optimizing under time-based vs. utilization-based attribution yields equivalent provisioning decisions (Appendix H).
@@ 3. Online Approach
Based on these findings, we propose a multi-horizon optimization which periodically updates forecasts and determines an optimal (or close-to-optimal) deployment and load balancing for every interval $i$, described in Algorithm 1.
Algorithm 1 Multi-Horizon Optimization
1:for $\alpha\leftarrow 0$ to I1I-1 do
2: # Long-term optimization
3:if $\alpha\equiv 0\pmod{\tau}$ then
4:  $(R^{i},C^{i})_{i=\alpha}^{I-1}\leftarrow$ update forecasts
6: # Short-term optimization
[…]
8: # Progress in time
[…]
We perform the optimization in two distinct steps:
@@
3.1. Automatic QoR Adaptation under an Annual Budget
The optimization problem in Section 3 minimizes emissions under a given $\text{QoR}_{\text{target}}$. However, especially for best-effort users, an alternative objective is to provide the highest possible service quality while staying within an annual carbon budget $B$. This setting is relevant when operators (or tenants) face internal sustainability targets, reporting requirements, or carbon pricing that effectively turns emissions into a constrained resource.
We model this by making $\text{QoR}_{\text{target}}$ a decision variable and optimizing
(8) $\max\quad$ $\text{QoR}_{\text{target}}$
@@ 3.1. Automatic QoR Adaptation under an Annual Budget
(9) $\sum_{i=\alpha}^{\omega}E^{i}\leq B$
(10) $\text{QoR}_{\text{target}}\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}.$
In practice, when re-optimizing at time $\alpha$, a part of the budget has already been spent. Let $E^{0},\dots,E^{\alpha-1}$ be the (observed) emissions of past intervals. Then the remaining budget is
(11) $B_{\text{rem}}(\alpha)=B-\sum_{i=0}^{\alpha-1}E^{i},$
and the long-term optimization step is solved over the remaining horizon with $\sum_{i=\alpha}^{\omega}E^{i}\leq B_{\text{rem}}(\alpha)$ while keeping all past decisions fixed. In the multi-horizon optimization (Algorithm 1), we use (8) as the objective for the long-term optimization step. This enables periodic updates of $\text{QoR}_{\text{target}}$ based on forecast updates and already-incurred emissions, while the short-term step continues to refine per-interval decisions within a smaller horizon.
@@ 4. Experiments
Scenario.  We investigate an LLM inference scenario, where a single user group is served by a LLaMA 3.1 (Dubey et al., 2024) model through the inference and serving engine vLLM (Kwon et al., 2023). Users can be fully served by LLaMA 3.1 8B ($\text{QoR}=0$), LLaMA 3.1 70B ($\text{QoR}=1$), or a fraction of both models ($0<\text{QoR}<1$). We consider $\mathcal{M}=\{\text{EC2 p4d.24xlarge}\}$, currently the only instance type with high-performance GPUs available on AWS, and define $p_{m}^{\text{attr}}=3781.8\,$W and $C_{m}^{\text{emb}}=135.3$ gCO2 based on (Teads Engineering, 2024). We model the machine performance $K$ using recent benchmarks of vLLM on a EC2 p4d.24xlarge for inference, which state a throughput of 11.57 requests per second for LLaMA 3.1 8B and 5.05 requests per second for LLaMA 3.1 70B (vLLM Team, [n. d.]).
@@
Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
  • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
  • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
  • In Australia, we consider Queensland (AU-QLD)
All carbon intensity data was provided by ElectricityMaps (Maps, 2024). We normalized all time zones to align the time of day with the request traces. Our forecast methodology is explained in Appendix E.
@@ 4.2. Upper-Bound Potential
To assess the potential of carbon-aware QoR adaptation, we assume perfect forecasts, and solve Equation (4) for $\gamma=\{$8 hours, 1 day, 1 week, 1 month}\}. Shorter periods show limited potential, as ACI is reported hourly and tends to vary only slowly over time. Note, that long validity periods can lead to reduced service quality over extended periods, see Appendix G. We optimize the objective until it is within 0.1 % of the optimal solution or after 1 hour.
@@ 4.3. Potential under Realistic Conditions
Figure 4 shows the gap to the upper-bound performance. Across all experiments, savings under realistic conditions reached $82\pm 6$ % of the upper-bound potential. Notably, even in regions with highly unpredictable request patterns (Random and Cell B, D and F), the online optimization does not show a significant performance drop.
4.4. Automatic QoR Adaptation under an Annual Budget
Lastly, we evaluate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 3.1 on Wiki (en) in three exemplary regions. For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$ (Section 4.2). Consequently, under perfect forecasts, an optimal solution would maintain $\text{QoR}_{\text{target}}\approx 0.5$ throughout the year.
Baselines.  We compare our approach to two greedy baselines that use the same forecasts as our method:
  1. (1)
    Greedy (constant) divides the remaining annual budget into equal per-interval budgets and greedily selects the highest feasible per-interval quality given the current forecast. Any remaining or overspent budget is evenly distributed across the remaining intervals.
  2. (2)
    Greedy (weighted) follows the same approach but weights the per-interval budgets by forecasted load and carbon intensity. Weightings are updated whenever forecasts are updated.
Figure 5. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL) under an annual carbon budget. Automatic adaptation provides consistent QoR throughout the year.
Figure 5 shows the resulting daily QoR under realistic forecasts. Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year. Greedy (weighted) is more stable but can still drift toward extreme QoR values toward the end of the year as forecast errors accumulate. In contrast, periodic long-term re-optimizations allow our approach to adjust $\text{QoR}_{\text{target}}$ over time, yielding more consistent QoR while staying within the annual budget.
@@ Appendix G Periods of Low Service Quality
Appendix H Equivalence of Provisioning Under Time- vs. Utilization-based Power Attribution
We prove that the optimal provisioning decision is the same under (i) utilization-based power attribution (3), where per-machine power depends on resource utilization, and (ii) time-based power attribution, where each active machine is assigned a utilization-independent attributed power draw. This is important, as utilization-based power attribution introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$. This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve.
Assumption 1.
$p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$. This reflects the basic property that active machines consume more power under load than when idle or turned off.
Assumption 2.
$n\leq 1$, i.e. the power model is concave, which is a common assumption for computing hardware as servers run more efficiently under high load (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007).
Assumption 3.
$p_{m_{1}}^{\text{idle}}<p_{m_{2}}^{\text{idle}}\ \Rightarrow\ p_{m_{1},q}^{\text{max}}<p_{m_{2},q}^{\text{max}}\quad\forall m_{1},m_{2}\in\widehat{\mathcal{M}}$. If a quality tier can be deployed on different machine types $\widehat{\mathcal{M}}$, we assume a strict ordering of these machines based on their idle and max power usage.
Theorem 1.
Let Assumptions 12, and 3 hold. When minimizing (4) subject to (5) – (7), the utilization-based power attribution,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl)\cdot\bigl(\text{util}^{i}_{q}\bigl)^{n},$
can be simplified to the same form as the time-based attribution
$\tilde{p}^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
Lemma 0.
Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines of the same type than strictly necessary. Formally, if Assumptions 1 and 2 hold, utilization-based attribution $f$ is strictly subadditive
(12) $f(x+y)<f(x)+f(y),$
where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
Proof of Lemma 2.
We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
Case 1: Linear power model ($n=1$):
(13) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
(14) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
(15) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
(16) $f(x+y)$ $<f(x)+f(y)$
Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $(0)\geq 0, ff$ is strictly subadditive on $(0,\infty)$.
(17) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
(18) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
(19) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl(}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl(}}_{>1}<0$
(20) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
Additionally, $f(0)=p_{m}^{\text{idle}}>0$. ∎
Proof of Theorem 1.
For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
[…]
We show, that both power attribution models yield $\check{D}^{i}$:
  • Feasibility with fewer machines is impossible: If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate (6).
  • Using extra machines is suboptimal for time-based power attribution: For $\tilde{p}^{i}_{m,q}$, any extra machine increases the objective by $p_{m,q}^{\text{max}}$.
  • Using extra machines is suboptimal for utilization-based power attribution: By Lemma 2, for $p^{i}_{m,q}$ there is no incentive to deploy extra machines of the same type. If a quality tier $q$ can be served by multiple machine types $\widehat{\mathcal{M}}$, under Assumption 3 (a strict ordering of these machines based on their idle and max power usage), if a more efficient machine type $m_{1}$ is still able to carry extra load, using a less efficient type $m_{2}$ instead can only increase the objective. Consequently, also across heterogeneous machine types, there is no incentive to add extra machines.
In either case (homogeneous or strictly ordered heterogeneous machines), any optimal solution must use exactly the minimal number of machines of each machine type. Because both the time-based and the utilization-based models thus select the same minimal deployment $\check{D}^{i}$, they yield equivalent optimal deployments. ∎